OpenStudy (anonymous):

Solve the system of linear equations below. 4x - 3y = 17 5x + 4y = 60 A. x = 2, y = -3 B. x = 5, y = 8 C. x = 8, y = 5 D. x = -8, y = -5

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

is it C?!

OpenStudy (hoblos):

4x - 3y = 17 *4 5x + 4y = 60 *3 16x - 12y = 68 15x + 12y = 180 add 31x = 248 x=8 y=5 C